How many pounds of coffee at 4.00 a pound must be mixed with coffee at 20.00 a pound to make 50 pounds of coffee that are worth 10.00 a pound?
Answer by checkley71(8403)

4X+20(50-X)=10*50
4X+1000-20X=500
-16X=500-1000
-16X=-500
X=-500/-16
X=31.25 POUNDS OF $4.00 COFFEE IS USED.
50-31.25=18.75 POUNDS OF $20.00 COFFEE IS USED.
PROOF:
4*31.25+20*18.75=500
125+375=500
500=500